# The Human Labour of Data Work: Capturing Cultural Diversity through World Wide Dishes

### v2

@@ 1. Introduction
Large-scale Machine Learning (ML) models underpin many essential digital services we interact with on an everyday basis, such as search engines, movie captions, and image generation tools. These ML models are socio-technical systems that both shape and are shaped by broader societal structures (Star, 1990; Pinch and Bijker, 1984). As a result, ML models inherit, reproduce, and amplify the values, norms, and biases of the societal structures in which they are developed. These models are imbued with social values through the datasets they are trained on and evaluated against (Birhane and Prabhu, 2021; Schuhmann et al., 2022). These datasets are typically constructed by scraping available data from the Internet (Cartografias da Internet, 2025; Mwema and Birhane, 2025; Podcast, 2025). The Internet, however, is not a neutral repository of human knowledge: it disproportionately represents perspectives from the Global North and reflects existing social, political, and economic power hierarchies (Luccioni et al., 2024; Hong et al., 2023). ML models built using web-scraped datasets tend to perpetuate harm against communities that are under-represented in training and evaluation data, creating a cycle that reinforces existing biases.
@@
To address representational gaps in training and evaluation data, ML researchers are actively exploring participatory approaches that directly engage marginalised communities in dataset development (Singh et al., 2024; Kirk et al., 2024; Ramaswamy et al., 2023). These efforts aim to address the “long-tailed” (Massiceti et al., 2021) distribution problem, where minority perspectives are vastly outnumbered relative to the overall distribution of data, by engaging historically marginalised communities in dataset development. Building on recent advances in community-engaged dataset development (Singh et al., 2024; Massiceti et al., 2021), our work illuminates the socio-technical infrastructure necessary to support community-engaged data curation approaches. We draw on CSCW scholarship investigating the labour underpinning large-scale systems and shift the focus from the dataset as the final product to the data work involved in its creation (Sambasivan et al., 2021; Scheuerman et al., 2021).
@@
We shift focus to the data work involved in building a dataset alongside community members by asking: How can researchers collect cultural data using a bottom-up, community-led approach? To answer our research question, we engage in a design retrospective (Irani and Silberman, 2016; Borowski et al., 2023) using first-person methods (Erete et al., 2021; Homewood, 2023; Wirfs-Brock et al., 2021; Cifor and Garcia, 2019; Desjardins et al., 2021; Howell et al., 2021) of our process of building the World Wide Dishes (WWD) dataset, a text and image dataset of culinary dishes and their associated cultural customs. The WWD dataset emerged from an effort to combat the “long-tailed” nature of datasets where cultural artefacts from marginalised groups are disproportionately under-represented due to myriad technical (Hong et al., 2024) and systemic causes (Noble, 2018). WWD collects data about food due to the cultural significance of cuisine: food is deeply intertwined with history, geography, and religious symbolism (UNESCO, 2003). Through our design retrospective, we surface how the labour performed by participatory mediators, whom we call Community Ambassadors (CAs) (Singh et al., 2024) in the context of WWD, facilitate bottom-up, community-led data collection efforts. We find that CAs make bottom-up, community-led data collection possible by doing three key things: building trust with community members, making participation accessible, and contextualising community values to support meaningful data collection. We argue that this labour is necessary to make participatory AI efforts successful in action.
@@
To ground our design retrospective of WWD, we first provide historical context that motivated the development of the WWD project in Section 3. We then describe the WWD dataset, the technical infrastructure built to facilitate data collection, and the stakeholders involved throughout the project in Section 4. In Section 5, we describe how we reflected on the process of developing and deploying WWD and unpack how our positionality impacts both the creation and analysis of our reflections. Through our retrospective analysis, we surface how CAs make bottom-up community-led approaches to dataset collection possible. In Section 6, we describe three ways in which CAs perform labour to support the development and deployment of WWD. In Section 7, we provide implications for methods and implementation of future participatory AI efforts, and reflect on how tensions that emerged during the data work were manifestations of deeper structural issues within the AI/ML development pipeline. We conclude our paper with a call for future work to invest in and build the necessary infrastructure to support participatory AI efforts in Section 8.
[…]
@@ 2.1. Bias in machine learning datasets
In this paper, we define bias as a reflection of human subjectivity (Miceli et al., 2022; Scheuerman et al., 2021). Data are produced by humans and therefore inextricable from human subjectivity (D’ignazio and Klein, 2023; Haraway, 2013). When a dataset is created, people make decisions about what deserves to be captured and how to classify what they are capturing (Bowker, 2000). Depending on who does the classifying and why the classification project is happening in the first place, some people may be accounted for while others are dismissed, invisible in the final dataset (Durkheim and Mauss, 2009). We say a dataset is biased when one perspective is disproportionately represented to the detriment of others as a result of these subjective decisions. In this paper, we follow calls by Miceli et al. (2022) and Barabas et al. (2020) to use bias as an analytical lens for examining datasets commonly used in ML development. This approach allows us to trace bias in ML model outputs back to the subjectivities embedded in datasets and, eventually, questions of who was involved in dataset creation (Barabas et al., 2020; Miceli et al., 2022).
@@
Datasets, carrying the worldviews and beliefs of their creators, can precipitate real-world harms when ML models trained on them reproduce these world views in their outputs (Blodgett et al., 2021; Birhane et al., 2021; Weidinger et al., 2023; Shankar et al., 2017; Scheuerman et al., 2019; Hall et al., 2024). For example, Buolamwini and Gebru (2018) showed that common facial recognition models often failed to “see” Black faces, which could result in a wide variety of real-world harms when facial recognition systems are used to verify workers’ identities (Watkins, 2023) or surveil low-income housing residents (MacMillan, 2023)Buolamwini and Gebru (2018) traced the roots of the failure of a facial recognition system to identify the faces of Black people back to training datasets lacking adequate and appropriate representation of Black people. Studying these representational lapses reveals the social values and power dynamics that shaped dataset construction, resulting in the systematic exclusion of Black people’s faces in digital datasets (Barabas et al., 2020).
@@
Machine learning research has extensively studied harmful model behaviours as they relate to gender (Berg et al., 2022; Hall et al., 2024; Rudinger et al., 2018; Birhane et al., 2021) and race (Currie et al., 2024; West et al., 2024; Buolamwini and Gebru, 2018). We respond to calls to expand the investigation of model breakdowns to additional axes (Weidinger et al., 2023) such as cultural representation. In this work, we use the term “culture” to mean “cultural heritage”: both tangible objects (e.g., monuments, sites) and intangible practices (e.g., traditions, rituals) that have a symbolic relationship to cultural identity  (Adilazuarda et al., 2024; Blake, 2000). We build on Miceli et al. (2022) and Barabas et al. (2020) to study breakdowns in ML model performance in the context of cultural representation through food. We choose food as a lens into culture because food is a salient cultural artefact. Food is considered to be a part of the intangible cultural heritage of a group that is shaped by a region’s history, geography, and even religious symbolism.  [^1] Intangible cultural heritage is considered a “mainspring of cultural diversity” and the social practices and rituals around food are so deeply intertwined with culture that they are recognised as “intangible cultural heritage” under UNESCO’s Convention for the Safeguarding of the Intangible Cultural Heritage. https://ich.unesco.org/en/convention. This is the primary international legal instrument in the field. Not all cultures, and therefore not all foods, are equally likely to be represented in datasets used in ML applications due to how, and by whom, these datasets are constructed. We trace absent and flattened cultural representation in model performance back to questions about who participates in dataset construction (see Section 3).
@@ 2.2. Mitigating bias in machine learning through crowdsourcing representative datasets
Social computing scholars have identified that datasets can encode biases which ML systems then reproduce (Gebru et al., 2021; Buolamwini and Gebru, 2018; Sweeney, 2013; Scheuerman et al., 2021; Scheuerman and Brubaker, 2024; Kapania et al., 2023; Sambasivan et al., 2021). In response to this issue, researchers have begun exploring ways to mitigate bias by crowdsourcing diverse values and expanding the range of worldviews represented in datasets (Singh et al., 2024; Romero et al., 2024). Crowdsourcing knowledge tends to take one of two forms: a top-down model that typically involves a corporate intermediary (e.g., ScaleAI, Amazon Mechanical Turk) that contracts community members as data workers and directs their work, and a bottom-up model that directly engages community members in the data collection and does not necessarily rely on a corporate intermediary to contract and oversee community members as data workers (Singh et al., 2024).
@@ 2.3. Bottom-up participatory opportunities for mitigating bias in machine learning
Whereas related work has commonly centred attention on top-down crowdsourcing, we focus our attention on bottom-up approaches. Our paper contributes to this discourse by extending the scope and opportunities for participation for mitigating bias in ML. Compared to top-down crowdsourcing, bottom-up and community-based crowdsourcing broaden the range of people who can contribute and can empower more people not only to share cultural artefacts, knowledge, and expertise, but also to help shape the parameters of their own participation (Denton et al., 2021; Delgado et al., 2023; Birhane et al., 2022). We characterise bottom-up, community-based crowdsourcing as efforts that allow for any community member to engage (e.g., they do not need to be formally employed as a data crowd worker), make space for community contributors to actively shape parts of the research process (e.g., they may take part in decisions about what kinds of data should be collected), and tend to—in large part—be volunteer-based.
[…]
While community-led efforts, such as the aforementioned initiatives, make significant strides in producing datasets with diverse community-based input, both research efforts foreground the dataset as the primary artefact of interest. In our contribution, we aim to build upon the precedent set by bottom-up crowdsourcing efforts by calling attention to the processes and complexities—such as designing accessible infrastructure to support data collection and working with communities to build trust and rapport—throughout the data collection process by which these types of datasets are created. We demonstrate how our effort, World Wide Dishes, fulfilled our characterisation of bottom-up community-based crowdsourcing by building the infrastructure to support any community member to engage with the project and actively shape the parameters of data collection in Section 4.2.
@@ 2.4. The pitfalls and potentials of participation
CSCW has a long and rich history of leveraging community-led participatory research (CBPR) methods to center socially marginalised groups’ epistemologies in the design of technology (Le Dantec and Fox, 2015; Sum et al., 2023; Bratteteig and Wagner, 2016; Harrington et al., 2019; Bannon et al., 2018; Pierre et al., 2021). CBPR is used as a method to mitigate the power imbalances between researchers and communities by empowering community members to shape the trajectory and design of research projects (Reason and Bradbury, 2008). For example, Asaro (2000) documents how, in light of impending technological changes in the workplace, workers themselves were invited to participate in the design of those technologies and thus had the power to shape the impacts of these technologies on their well-being. Harrington et al. (2019) considers the role of research setting on community access, and therefore the accessibility of participation for targeted groups. In CBPR, practitioners strive to overcome their status as outsiders by building trust through affective and moral connections with community members (Le Dantec and Fox, 2015; McMillan and Chavis, 1986). Taken together, CBPR literature overwhelmingly stresses the need for researchers to become imbricated with the communities they are studying; through building these relationships, they become invested in the outcomes of the research from an insider, community member perspective.
@@
While participatory methods have the potential to give communities a greater say over the design of technology, many of these efforts fall short in meaningfully shifting power from the researchers to the researched (Birhane et al., 2021; Ojewale et al., 2024; Young et al., 2024). Participation takes work on behalf of communities and the process of participating can be negatively impactful for participants. For example, Epstein (2008) critiques participatory design efforts for their tendency to “other” socially marginalised groups in the course of trying to single them out for inclusion in research. Pierre et al. (2021) documents how communities assume the “epistemic burden” of producing data about their lived experiences in participatory research efforts. Puri and and (2007) expand upon how participatory design efforts in rural communities require careful attention to building support structures that lower the barrier to entry and build intrinsic motivation within community members to engage in participation. Building on (Puri and and, 2007), we call attention to the role of participatory mediators in building pathways for equitable community participation in dataset development. In CSCW literature, mediators are shown to be a crucial component of technical system adoption in organisations (Okamura et al., 1995; Friedman and Cornford, 1989; Grudin, 1988). We extend this concept of mediators to a participatory design project, showing how community ambassadors can be used to mediate interactions between community members and the broader research project throughout its development and deployment phase. We argue that participatory mediators are a crucial component of participatory efforts, especially when they are geared towards including people from socially marginalised communities, and propose design recommendations for implementing future participatory AI projects.

### v3

@@ Abstract.
Data work, cultural representation, positionality, community-based design
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW492publicationmonth: 11doi: 10.1145/3757673ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Computer visionccs: Social and professional topics Cultural characteristics
@@ 1. Introduction
Large-scale Machine Learning (ML) models underpin many essential digital services we interact with on an everyday basis, such as search engines, movie captions, and image generation tools. These ML models are socio-technical systems that both shape and are shaped by broader societal structures (Star, 1990; Pinch and Bijker, 1984). As a result, ML models inherit, reproduce, and amplify the values, norms, and biases of the societal structures in which they are developed. These models are imbued with social values through the datasets they are trained on and evaluated against (Birhane and Prabhu, 2021; Schuhmann et al., 2022). These datasets are typically constructed by scraping available data from the Internet (Cartografias da Internet, 2025; Mwema and Birhane, 2025; Podcast, 2025). The Internet, however, is not a neutral repository of human knowledge: it disproportionately represents perspectives from the Global North and reflects existing social, political, and economic power hierarchies (Luccioni et al., 2024; Hong et al., 2023). ML models built using web-scraped datasets tend to perpetuate harm against communities that are underrepresented in training and evaluation data, creating a cycle that reinforces existing biases (Magomere et al., 2025; Katzman et al., 2023; Qadri et al., 2025).
@@
To address representational gaps in training and evaluation data, ML researchers are actively exploring participatory approaches that directly engage marginalised communities in dataset development(Singh et al., 2024; Kirk et al., 2024; Ramaswamy et al., 2023). These efforts aim to address the “long-tailed”(Massiceti et al., 2021) distribution problem—where minority perspectives are vastly outnumbered relative to the overall distribution of data—by engaging historically marginalised communities in dataset development. Building on recent advances in community-engaged dataset development (Singh et al., 2024; Massiceti et al., 2021), our work illuminates the socio-technical infrastructure necessary to support community-engaged data curation approaches. We draw on CSCW scholarship investigating the labour underpinning large-scale systems and shift focus from the dataset as the final product to the data work involved in its creation (Sambasivan et al., 2021; Scheuerman et al., 2021).
@@
We shift focus to the data work involved in building a dataset alongside community members by asking: How can researchers collect cultural data using a bottom-up, community-led approach? To answer our research question, we use first-person methods (Erete et al., 2021; Homewood, 2023; Wirfs-Brock et al., 2021; Cifor and Garcia, 2019; Desjardins et al., 2021; Howell et al., 2021) to engage in a design retrospective (Irani and Silberman, 2016; Borowski et al., 2023) of our process of building the World Wide Dishes (WWD) dataset, a text and image dataset of culinary dishes and their associated cultural customs. The WWD dataset emerged from an effort to combat the “long-tailed” nature of datasets where cultural artefacts from marginalised groups are disproportionately underrepresented due to myriad technical (Hong et al., 2024) and systemic causes (Noble, 2018). WWD collects data about food due to the cultural significance of cuisine: food is deeply intertwined with history, geography, and religious symbolism (UNESCO, 2003). Through our design retrospective, we surface how the labour performed by participatory mediators, whom we here call Community Ambassadors (CAs) (Singh et al., 2024), facilitates bottom-up, community-led data collection efforts. We find that CAs make bottom-up, community-led data collection possible by doing three key things: building trust with community members, making participation accessible, and contextualising community values to support meaningful data collection. We argue that this labour is necessary to make participatory AI efforts successful in action.
@@
To ground our design retrospective of WWD, we first provide historical context that motivated the development of the WWD project in Section 3. We then describe the WWD dataset, the technical infrastructure built to facilitate data collection, and the stakeholders involved throughout the project in Section 4. In Section 5, we describe how we reflected on the process of developing and deploying WWD and unpack how our positionality impacts both the creation and analysis of our reflections. Through our retrospective analysis, we surface how CAs make bottom-up, community-led approaches to dataset collection possible. In Section 6, we describe three ways in which CAs perform labour to support the development and deployment of WWD. In Section 7, we provide implications for methods and implementation of future participatory AI efforts, and reflect on how tensions that emerged during the data work were manifestations of deeper structural issues within the AI/ML development pipeline. We conclude our paper with a call for future work to invest in and build the necessary infrastructure to support participatory AI efforts in Section 8.
We build on emergent efforts in the participatory AI space that develop ML models by community members, for community members, such as Masakhane ($\forall$ et al., 2020b) and Cohere’s AYA Project(Singh et al., 2024). We extend the participatory design literature in CSCW (Irani and Silberman, 2016; Howell et al., 2021; Le Dantec and Fox, 2015; Harrington et al., 2019) by characterising how a participatory AI effort unfolds in practice and making visible the labour performed by participatory mediators throughout the research process. Our paper shows how bottom-up, community-driven ML development efforts are contingent on labour performed by CAs. These participatory mediators are an essential infrastructural component that makes participatory design projects viable. We present suggestions for how future participatory AI efforts can account for and support the labour of participatory mediators.
@@ 2.1. Bias in machine learning datasets
In this paper, we define bias as a reflection of human subjectivity (Miceli et al., 2022; Scheuerman et al., 2021). Data are produced by humans and therefore inextricable from human subjectivity (D’ignazio and Klein, 2023; Haraway, 2013). When a dataset is created, people make decisions about what deserves to be captured and how to classify what they are capturing (Bowker, 2000). Depending on who does the classifying and why the classification project is being undertaken, some people may be accounted for while others are dismissed, invisible in the final dataset (Durkheim and Mauss, 2009). We say a dataset is biased when one perspective is disproportionately represented to the detriment of others as a result of these subjective decisions. In this paper, we follow calls by Miceli et al. (2022) and Barabas et al. (2020) to use bias as an analytical lens for examining datasets commonly used in ML development. This approach allows us to trace bias in ML model outputs back to the subjectivities embedded in datasets and, eventually, questions of who was involved in dataset creation (Barabas et al., 2020; Miceli et al., 2022).
@@
Datasets, carrying the worldviews and beliefs of their creators, can precipitate real-world harms when ML models trained on them reproduce these worldviews in their outputs (Blo, [n. d.]; Birhane et al., 2021; Weidinger et al., 2023; Shankar et al., 2017; Scheuerman et al., 2019; Hall et al., 2024). For example, Buolamwini and Gebru (2018) showed that common facial recognition models often failed to “see” Black faces, which could result in a wide variety of real-world harms when facial recognition systems are used to verify workers’ identities (Watkins, 2023) or surveil low-income housing residents (MacMillan, 2023). Buolamwini and Gebru (2018) traced the roots of the systems’ failures back to the training datasets having lacked adequate and appropriate representation of Black people. Studying these representational lapses reveals the social values and power dynamics that shaped dataset construction, resulting in the systematic exclusion of Black people’s faces in digital datasets (Barabas et al., 2020).
@@
Machine learning research has extensively studied harmful model behaviours as they relate to gender (Berg et al., 2022; Hall et al., 2024; Rudinger et al., 2018; Birhane et al., 2021) and race (Currie et al., 2024; West et al., 2024; Buolamwini and Gebru, 2018). We respond to calls to expand the investigation of model breakdowns to additional axes (Weidinger et al., 2023) such as cultural representation. In this work, we use the term “culture” to mean “cultural heritage”: both tangible objects (e.g., monuments, sites) and intangible practices (e.g., traditions, rituals) that have a symbolic relationship to cultural identity  (Adilazuarda et al., 2024; Blake, 2000). We build on Miceli et al. (2022) and Barabas et al. (2020) to study breakdowns in ML model performance in the context of cultural representation through food. We choose food as a lens into culture because food is a salient cultural artefact. Food is a part of the intangible cultural heritage of a group that is shaped by a region’s history, geography, and even religious symbolism. [^1] Intangible cultural heritage is considered a “mainspring of cultural diversity” and the social practices and rituals around food are so deeply intertwined with culture that they are recognised as “intangible cultural heritage” under UNESCO’s Convention for the Safeguarding of the Intangible Cultural Heritage. https://ich.unesco.org/en/convention. This is the primary international legal instrument in the field. Not all cultures, and therefore not all foods, are equally likely to be represented in datasets used in ML applications due to how and by whom these datasets are constructed. We trace absent and flattened cultural representation in model performance back to questions about who participates in dataset construction (see Section 3).
@@ 2.2. Mitigating bias in machine learning through crowdsourcing representative datasets
Social computing scholars have identified that datasets can encode biases which ML systems then reproduce (Gebru et al., 2021; Buolamwini and Gebru, 2018; Sweeney, 2013; Scheuerman et al., 2021; Scheuerman and Brubaker, 2024; Kapania et al., 2023; Sambasivan et al., 2021). In response to this issue, researchers have begun exploring ways to mitigate bias by crowdsourcing diverse values and expanding the range of worldviews represented in datasets (Singh et al., 2024; Romero et al., 2024). Crowdsourcing knowledge tends to take one of two forms: a top-down model that typically involves a corporate intermediary (e.g., ScaleAI, Amazon Mechanical Turk) that contracts community members as data workers and directs their work, or a bottom-up model that directly engages community members in the data collection and does not necessarily rely on a corporate intermediary to contract and oversee community members as data workers (Singh et al., 2024).
@@ 2.3. Bottom-up participatory opportunities for mitigating bias in machine learning
Whereas related work has commonly centred attention on top-down crowdsourcing, we focus our attention on bottom-up approaches. Our paper contributes to this discourse by extending the scope and opportunities for participatory methods supporting bias mitigation in ML. Compared to top-down crowdsourcing, bottom-up and community-based crowdsourcing broaden the range of people who can contribute and can empower more people not only to share cultural artefacts, knowledge, and expertise, but also to help shape the parameters of their own participation (Denton et al., 2021; Delgado et al., 2023; Birhane et al., 2022). We characterise bottom-up, community-based crowdsourcing as efforts that allow for any community member to engage (e.g., they do not need to be formally employed as a data crowd worker), make space for community contributors to actively shape parts of the research process (e.g., they may take part in decisions about what kinds of data should be collected), and tend, in large part, to be volunteer-based.
While top-down methods for crowdsourcing might be easier to scale, recent efforts to build large-scale datasets for ML applications using bottom-up, community-based crowdsourcing techniques have shown great promise. For example, in 2020, $\forall$ et al. (2020b) introduced Masakhane, which is an ongoing open-source, continent-wide online research effort for machine translation for African languages. Masakhane has also consistently demonstrated ethical participatory research design for machine translation and natural language processing in Africa ($\forall$ et al., 2020a; Adelani et al., 2022; ogu, [n. d.]). Cohere’s AYA dataset (Singh et al., 2024) also presents a framework for large-scale community data collection. Their collaborators spanned 119 countries and were able to create a multilingual text dataset comprising 513 million instances across 114 languages. All their systems have been open-sourced. The Masakhane and AYA Project initiatives demonstrate that bottom-up, community-based crowdsourcing techniques can yield large-scale datasets that are grounded in community expertise and empower everyday community members to participate in dataset construction without having to be formal data workers.
While community-led efforts, such as the aforementioned initiatives, make significant strides in producing datasets with diverse community-based input, both research efforts foreground the dataset as the primary artefact of interest. In our contribution, we aim to build upon the precedent set by bottom-up crowdsourcing efforts by calling attention to the processes and complexities throughout the data collection process by which these types of datasets are created—such as by designing accessible infrastructure to support data collection and working with communities to build trust and rapport. We demonstrate how our effort, World Wide Dishes, fulfilled our characterisation of bottom-up community-based crowdsourcing by building the infrastructure to support any community member to engage with the project and actively shape the parameters of data collection in Section 4.2.
@@ 2.4. The pitfalls and potentials of participation
CSCW has a long and rich history of leveraging community-based participatory research (CBPR) methods to centre socially marginalised groups’ epistemologies in the design of technology (Le Dantec and Fox, 2015; Sum et al., 2023; Bratteteig and Wagner, 2016; Harrington et al., 2019; Bannon et al., 2018; Pierre et al., 2021). CBPR is used as a method to mitigate the power imbalances between researchers and communities by empowering community members to shape the trajectory and design of research projects (Reason and Bradbury, 2008). For example, Asaro (2000) documents how, in light of impending technological changes in the workplace, workers themselves were invited to participate in the design of those technologies and thus had the power to shape the impacts of these technologies on their wellbeing. Harrington et al. (2019) consider the role of research setting on community access, and therefore the accessibility of participation for targeted groups. In CBPR, practitioners strive to overcome their status as outsiders by building trust through affective and moral connections with community members (Le Dantec and Fox, 2015; McMillan and Chavis, 1986). Taken together, CBPR literature overwhelmingly stresses the need for researchers to become imbricated with the communities they are studying; through building these relationships, they become invested in the outcomes of the research from an insider, community member perspective.
@@
While participatory methods have the potential to give communities a greater say over the design of technology, many of these efforts fall short in meaningfully shifting power from the researchers to the researched (Birhane et al., 2021; Ojewale et al., 2024; Young et al., 2024). Participation takes work on behalf of communities and the process of participating can be negatively impactful for participants. For example, Epstein (2008) critiques participatory design efforts for their tendency to “other” socially marginalised groups in the course of trying to single them out for inclusion in research. Pierre et al. (2021) document how communities assume the “epistemic burden” of producing data about their lived experiences in participatory research efforts. Puri and Sahay (2007) expand upon how participatory design efforts in rural communities require careful attention to building support structures that lower the barrier to entry and build intrinsic motivation within community members to engage in participation. Building on (Puri and Sahay, 2007), we call attention to the role of participatory mediators in building pathways for equitable community participation in dataset development. In CSCW literature, mediators are shown to be a crucial component of technical system adoption in organisations (Okamura et al., 1995; Friedman and Cornford, 1989; Grudin, 1988). We extend this concept of mediators to a participatory design project, showing how community ambassadors can be used to mediate interactions between community members and the broader research project throughout its development and deployment phase. We argue that participatory mediators are a crucial component of participatory efforts, especially when those efforts are geared towards including people from socially marginalised communities, and propose design recommendations for implementing future participatory AI projects.